OpenStudy (calculator):

Confusing stoichiometry question, 0.144 g of an aluminium compound X react with an excess of water, to produce a gas. This gas burns completely in O2 to form H2O and 72 cm3 of CO2 only. The volume of CO2 was measured at room temperature and pressure. What could be the formula of X? [C = 12.0, Al = 27.0; 1 mole of any gas occupies 24 dm3 at room temperature and pressure]

OpenStudy (mos1635):

This gas burns completely in O2 to form H2O and 72 cm3 of CO2 only. so gas is CxHyOz 72 cm3=72*10^-3 dm3=72*10^-3 dm3/24 dm3=3*10^-3 moles CxHyOz +(4x+y-2z)/4 O2 --> xCO2 +y/2 H2O 1 moles gives x moles n moles n*x moles nx=3*10^-3 (1)............... are you sure there is no more info???????? it is most unlikly to continue with so meny variables......

OpenStudy (calculator):

Its a multi-choice A. Al2C3 B. Al3C4 C. Al4C3 D. Al5C3

OpenStudy (anonymous):

the answer is Al4C3.. this is a past as level question..Aluminium has 3 outer electrons and carbon has 4 ..so they interchange

OpenStudy (anonymous):

do you understand it?